OpenStudy (anonymous):

Please help I take an online class and the teacher gave us no notes can y'all explain how to do this? Given that the angle theta is in the first quadrant and given the point (12/13,5/13) is on the terminal side of the angle determine the exact values of the six trigonometric function of the real number t Sin Cos Tan Csc and Etc.

OpenStudy (anonymous):

every trig function can be composed of sin or cos, or both ... given some (x,y), r = sqrt(x^2+y^2) sin = y/r cos = x/r tan = sin/cos sec = 1/cos csc = 1/sin cot = 1/cot = cos/sin

OpenStudy (anonymous):

i beleive in this case, 5,12,13 is a pythag triple and so r=1

OpenStudy (anonymous):

ok so if sin=y/r than the answer is 5/13? or do we plug 5/13 into the calculator like sin(5/13) ??